What is the elasticity of the demand curve between points A and C? Round your answer to 2 decimal places.13 14 18 44 a
Answer #1
Price Quantity Demanded
$23 13
$9 18

Price elasticity of demand (Midpoint formula)

Price Elasticity of Demand = Percentage change in quantity demand Percentage change in price

Q2-01 Percentage Change in Quantity Demand = 7 (Q2 +01)/2

18 – 13 Percentage Change in Quantity Demand =; (18+ 13)/2

Percentage Change in Quantity Demand = 32.25

P2 - PI Percentage Change in Price = (P2 + PL/* 100

9 – 23 Percentage Change in Price = (9 +23)/0 * 100

Percentage change in Price = -87.5

Price Elasticity of Demand = Percentage change in quantity demand Percentage change in price

32.25 Price Elasticity of Demand = 3 -87.5

Price Elasticity of Demand = -0.36

Negative sign is ignored in price elasticity of demand because of the inverse relationship between price and quantity demand

Price Elasticity of Demand = 0.36